# OntoURL: A Benchmark for Evaluating Large Language Models on Symbolic Ontological Understanding, Reasoning and Learning

### v2

@@ 3.3 Task Definition
Based on the taxonomy in Section 3.1 and the pipeline in Section 3.2, we developed a series of tasks corresponding to the three ontological capabilities. An overview of these tasks is provided in Table 1, and detailed descriptions along with specific examples are presented in Appendix D.
For Understanding capability, tasks evaluate a model’s ability to comprehend explicitly defined ontological elements, including class definitions (U1), class relationships (U2), property domains (U3), instance classifications (U4), and instance definitions (U5).
The Reasoning capability increases complexity by requiring inference over implicit knowledge not explicitly presented in the ontology. Inferred relation reasoning (R1) extends task U2 by shifting from explicit to inferred class relationships. Similarly, constraint reasoning (R2) and instance class reasoning (R3) are inference-based counterparts of tasks U3 and U4, respectively. Tasks R4 and R5 introduce more advanced logical inference: SWRL-based logic reasoning (R4) involves reasoning over rules defined in the Semantic Web Rule Language (horrocks2004swrl, , SWRL), encompassing conjunctions, property chains, and conditional implications. Description logic reasoning (R5) focuses on reasoning with description logic, where models must interpret formal expressions and perform deductive inference over constructs such as $\exists$, $\forall$, $\sqcap$, and numerical restrictions (e.g., $\geq n$, $\leq m$).
@@
R5 description logic reasoning T/FQ Accuracy 2,560
Learning L1 class definition generation Generation ROUGE-L 2,936
@@ 3.4 Benchmark Statistics
The domain distribution (Figure 4b) is directly related to the quantity and scale of the collected ontologies. The Sciences domain represents the largest portion (28.4%), as it encompasses 8 ontologies, including extremely large resources like the Gene Ontology ashburner2000gene ; gene2023gene , Cell Ontology diehl2016cell , and ChEBI degtyarenko2007chebi . The Health & Medicine domain follows as the second largest (22.7%), while the Legal domain represents the smallest share (2.1%), comprising only four relatively small ontologies.
Figure 4: Question Distribution of OntoURL Tasks and Domains. Additional statistics, such as average lengths of questions, options, and answers, are provided in Appendix D.
@@ 4 Evaluation
4.1 Setup
We evaluate LLMs under both zero-shot and few-shot settings across all 15 tasks in OntoURL. For the zero-shot scenario, the input to the LLMs consists solely of task instructions, questions, and answer options (where applicable). In the few-shot setting, we provide two or four carefully selected examples for each task to demonstrate the expected reasoning pattern and output format. As shown in Table 1, we use task-appropriate metrics: Accuracy for multiple-choice and true/false questions (tasks U1-U5, R1-R5), ROUGE-L for text generation (task L1), and F1 score for structured outputs such as triples or tuples (tasks L2-L5). We apply regular expressions to extract valid triples or tuples from the model’s responses to mitigate the impact of irrelevant text. The hyperparameters and hardware configurations are detailed in Appendix F. All instructions and prompts are provided in Appendix D.
@@ 4.2 Evaluated Models
We evaluate a diverse set of 20 language models, which can be categorized into three groups. General-purpose LLMs include 14 widely used open-source models across various parameter scales: Qwen2.5-(3B, 7B, 14B, 32B, 72B) qwen2025qwen25technicalreport , QWQ-32B, Phi4-4B abdin2024phi4technicalreport , LLaMA3.1-8B, LLaMA3.3-70B grattafiori2024llama3herdmodels , Aya-Expanse-(8B, 32B), InternLM3-8B, Mistral-8B, and Mistral-small. Ontology-trained LLMs comprise two task-specialized models—Ollm-wiki and Ollm-arxiv e2eol2024lo —which are fine-tuned from Mistral-7B jiang2023mistral7b on Wikipedia category and arXiv taxonomy data, respectively. Domain-specialized LLMs include SaulLM-7B (legal domain) colombo2024saullm7bpioneeringlargelanguage , BioMistral-7B (sciences) labrak2024biomistralcollectionopensourcepretrained , OpenBioLLM-8B (biomedicine), and Finance-Chat-7B (finance) cheng2024adaptinglargelanguagemodels . These models are included to provide a complementary perspective by assessing how domain-specific pretraining affects performance on ontological tasks. Links to all model repositories are provided in Appendix F.
@@ 4.3 Experimental Results and Analysis
We present the performance of models in Table 2 under the zero-shot setting. Additional results for few-shot experiments are provided in AppendixG. In the following analysis, we discuss the results from three perspectives: model performance, ontological capabilities, and performance on specific domains. To have a better view of the outputs, we also include case studies in Appendix H.
Table 2: Main results (%) of 16 LLMs (grouped by size) across three capability levels under the zero-shot setting. ★ indicates the best-performing model overall, while ❖ denotes the best-performing model within its size category.
Model Understanding (Acc.) Reasoning (Acc.) Learning (Rouge-L, F1)
@@
Qwen2.5-7B ❖ 83.1 90.6 77.6 90.1 83.6 87.6 88.2 73.9 66.0 68.6 5.6 0.4 0.1 0.3 16.2
Ollm-wiki-7B 74.3 84.5 67.2 81.4 77.0 65.2 83.3 53.4 57.3 59.3 6.0 0.1 0.0 0.2 0.1
Ollm-arxiv-7B 74.1 84.4 67.5 81.5 77.0 64.2 82.8 53.1 56.6 58.4 6.0 0.1 0.0 0.0 8.3
LLaMA3.1-8B 79.8 87.4 74.9 88.4 81.1 79.8 84.2 72.3 62.2 68.9 6.4 0.1 0.0 0.1 15.3
Ministral-8B 78.9 88.8 62.4 83.9 79.5 81.0 88.4 60.1 62.4 52.7 10.3 0.1 0.0 0.1 16.4
Internlm3-8B 83.1 90.9 72.0 88.9 82.4 88.5 90.5 73.8 67.2 62.9 6.3 0.2 0.0 0.4 12.0
Aya-8B 77.1 85.8 62.4 83.8 77.9 73.0 78.0 62.6 57.4 63.6 7.5 0.1 0.0 0.0 6.3
@@
Qwen2.5-14B 86.6 92.0 75.5 91.4 85.8 89.6 94.0 76.4 71.2 63.6 5.6 0.1 0.1 1.0 19.5
Mistral-22B 83.9 90.4 69.6 88.6 84.4 86.3 86.9 69.3 64.0 54.3 7.5 0.1 0.0 0.8 15.8
Qwen2.5-32B ❖ 88.0 90.6 81.9 91.2 87.2 89.7 95.5 76.8 72.4 68.4 5.7 1.6 0.1 1.5 20.3
QwQ-32B 82.2 89.6 77.1 88.9 81.5 84.0 92.5 70.8 60.6 63.4 6.1 1.1 0.2 1.0 18.0
Aya-32B 81.2 90.5 61.6 89.7 82.3 85.5 83.1 70.3 66.0 68.8 6.5 0.1 0.1 0.5 19.3
@@
LLaMA3.3-70B 88.0 94.1 76.8 91.8 90.0 91.9 92.9 76.8 70.9 64.2 5.7 0.1 0.0 0.7 20.2
Qwen2.5-72B ★ 89.1 92.6 84.3 92.6 89.4 92.1 93.4 77.5 75.6 68.4 6.0 0.1 0.0 1.0 21.6
Model Performance and Impact of Model Scale
@@
The two largest models, LLaMA3.3-70B and Qwen2.5-72B, consistently achieve the best performance across tasks. Notably, the Qwen architecture shows robust results at all sizes, outperforming other architectures of comparable scale. In contrast, Ollm, which is specifically trained for ontology construction, performs relatively poorly—likely due to its specialization in hierarchical generation rather than general understanding or reasoning.
@@
Model scale correlates strongly with performance, especially in understanding and reasoning. For instance, Qwen2.5-72B achieves top scores of 92.6 on U4, 93.4 on R2, and 21.6 F1-score on L5. Similarly, LLaMA3.3-70B scores 91.8 on U4, 92.9 on R2, and 20.2 on L5. This pattern is even more pronounced within the same architecture: across Qwen, Mistral, Aya, and LLaMA, larger models consistently outperform smaller models on most tasks.
Ontological Understanding Capability of LLMs
@@
Tasks U1 to U5 demonstrate that LLMs generally perform well on ontological understanding, particularly in recognizing hierarchical structures. This is reflected in the high accuracy on U2 (class relations) and U4 (instance classification), with scores ranging from 80% to 94%. In contrast, performance is less consistent on definition- and property-based tasks. U1 (class definition), U3 (property domain), and U5 (instance definition) show notable gaps in certain models, for instance, Aya-8B scores only 77.1%, 62.4%, and 77.9% on these tasks, respectively. As suggested in the examples (Appendix D), U1, U3 and U5 require the knowledge of highly specialized domains and the distractors closely resemble the correct answer.
Ontological Reasoning Capability of LLMs
@@
Reasoning tasks present greater challenges for LLMs than understanding tasks. Task R1 (inferred relation reasoning) is a difficult variant of task U2 (class relation understanding), requiring inference to identify class relationships not explicitly defined. As expected, models generally perform 3-4 percentage points worse on R1 than on U2, with the most dramatic decrease observed in Ollm-arxiv-7B (from 84.4% to 64.2%). Similarly, R3 (instance class reasoning) functions as the reasoning-intensive counterpart to U4. Performance on R3 (60-70%) demonstrates a substantial decline compared to U4 (80-90%). These results indicate that when reasoning across multiple relationships is required, LLM performance deteriorates significantly.
@@
Tasks involving logical expressions, such as R4 (SWRL-based reasoning) and R5 (description logic reasoning), are even more difficult. Compared to natural language-based reasoning (R1, R2 and R3), model performance drops more dramatically when logical operators are involved, with scores ranging between 60% and 75%. Even the best-performing model, Qwen2.5-72B, achieves only 75.6% and 68.8% on these two tasks, respectively. This highlights a significant limitation of current LLMs: difficulty in executing precise symbolic reasoning over formally structured ontologies.
Ontological Learning Capabilities of LLMs
@@
Although direct comparisons is limited due to different evaluation metrics, generation tasks are shown to be more challenging, comparing with understanding and reasoning tasks. In L1 (class definition generation), models struggle to generate definitions, as evidenced by low ROUGE-L (typically below 10). This poor performance likely stems from the challenges of domain-specific definition generation, which requires not only describing the target class but also distinguishing it from adjacent concepts (e.g. its superclasses). Unlike human domain experts who possess comprehensive ontological perspectives, LLMs struggle to make such fine-grained semantic distinctions within specialized domains.
@@
Similar limitations appear in structural construction tasks. In L2 (class hierarchy construction), L3 (property relation construction), and L4 (constraint construction), models frequently fail to produce syntactically valid triples in zero-shot settings. Performance improves modestly in two- and four-shot settings (Appendix G), but output quality remains low, often featuring ill-defined or hallucinated relations (Appendix H). For L5 (ontology alignment), models achieve slightly higher F1 scores in the range of 10-20%. But for practical ontology applications, the performance of the evaluated LLMs on all learning tasks remains substantially poor for reliable deployment.
@@
Figure 5: The performance of Qwen2.5-7B, 72B (green and blue) and four domain-specific LLMs (yellow) across three capabilities for four domains.
In addition to the task-based evaluation, Figure 5 compares the performance of Qwen2.5-7B, Qwen2.5-72B, and four domain-specific LLMs (Finance-Chat-7B, OpenBioLLM-8B, BioMistral-7B, and SaulLM-7B) across four key domains. For the Sciences domain, we retained only the Biology-related tasks. To simplify the computation of the mean scores for the learning task, we omitted the Rouge-L scores of L1.
As expected, the trends observed earlier hold here as well: larger models consistently outperform smaller ones, and models struggle more on reasoning and learning tasks than on understanding tasks. Comparing across domains, performance in Legal and the Sciences (Biology) lags behind the other two domains, with the gap most pronounced in Biology—likely a reflection of the greater complexity and specialized knowledge required.

### v3

@@ 3.3 Task Definition
R5 description logic reasoning T/FQ Accuracy 882
Learning L1 class definition generation Generation BERTScore 2,936
@@
We developed a series of tasks corresponding to the three ontological capabilities. An overview of these tasks is provided in Table 1.
For Understanding capability, tasks evaluate a LLM’s ability to comprehend explicitly defined ontological elements, including class definitions (U1), class relationships (U2), property domains (U3), instance classifications (U4), and instance definitions (U5).
The Reasoning capability increases complexity by requiring inference over implicit knowledge not explicitly presented in the ontology. Inferred relation reasoning (R1) extends task U2 by shifting from explicit to inferred class relationships. Similarly, constraint reasoning (R2) and instance class reasoning (R3) are inference-based counterparts of tasks U3 and U4, respectively. Tasks R4 and R5 introduce more advanced logical inference: SWRL-based logic reasoning (R4) involves reasoning over rules defined in the Semantic Web Rule Language (Horrocks et al., 2004, SWRL), encompassing conjunctions, property chains, and conditional implications. Description logic reasoning (R5) focuses on reasoning with description logic, where models must interpret formal expressions and perform deductive inference over constructs such as $\exists$, $\forall$, $\sqcap$, and numerical restrictions (e.g., $\geq n$, $\leq m$).
@@ 3.4 Benchmark Statistics
Figure 4: Question distribution of OntoURL tasks and domains. Additional statistics, such as average lengths of questions, options, and answers, are provided in Appendix B.
@@
The domain distribution (Figure 4b) is directly related to the quantity and scale of the collected ontologies. The Sciences domain represents the largest portion (28.4%), as it encompasses 8 ontologies, including extremely large resources like Gene Ontology Ashburner et al. (2000); Aleksander et al. (2023) and Cell Ontology Diehl et al. (2016). The Health & Medicine domain follows as the second largest (22.7%), while the Legal domain represents the smallest share (2.1%), comprising only four relatively small ontologies.
@@ 4 Evaluation
We evaluate LLMs under both zero-shot and few-shot settings across all 15 tasks in OntoURL. For the zero-shot scenario, the input to the LLMs consists solely of task instructions, questions, and answer options (where applicable). In the few-shot setting, we provide two or four carefully selected examples for each task to demonstrate the expected reasoning pattern and output format. As shown in Table 1, we use task-appropriate metrics: Accuracy for multiple-choice and true/false questions (tasks U1-U5, R1-R5), BERTScore F1 Zhang et al. (2020) for text generation (task L1), and F1 score for structured outputs such as triples or tuples (tasks L2-L5). We apply regular expressions to extract valid triples or tuples from the model’s responses to mitigate the impact of irrelevant text. The hyperparameters and configurations are detailed in Appendix D.
@@ 4.1 Evaluated Models
We evaluate a diverse set of 20 language models, which can be categorized into three groups. General-purpose LLMs include 14 widely used open-source models across various parameter scales: Qwen2.5-(3B, 7B, 14B, 32B, 72B) Yang et al. (2025), QWQ-32B, Phi4-4B Abdin et al. (2024), LLaMA3.1-8B, LLaMA3.3-70B Grattafiori and et. al. (2024), Aya-Expanse-(8B, 32B), InternLM3-8B, Mistral-8B, and Mistral-small. Ontology-trained LLMs comprise two task-specialized models—Ollm-wiki and Ollm-arxiv Lo et al. (2024)—which are fine-tuned from Mistral-7B Jiang et al. (2023) on Wikipedia category and arXiv taxonomy data, respectively. Domain-specialized LLMs include SaulLM-7B (legal domain) Colombo et al. (2024), BioMistral-7B (sciences) Labrak et al. (2024), OpenBioLLM-8B (biomedicine), and Finance-Chat-7B (finance) Cheng et al. (2024). These models are included to provide a complementary perspective by assessing how domain-specific pretraining affects performance on ontological tasks. Links to all model repositories are provided in Appendix D.
@@ 4.2 Experimental Results and Analysis
We present the performance of models in Table 2 under the zero-shot setting. In the following analysis, we discuss the results from three perspectives: model performance, ontological capabilities, and performance on specific domains.
Model Understanding (Acc.) Reasoning (Acc.) Learning (BERTScore, F1)
@@
Qwen2.5-7B ❖ 83.1 90.6 77.6 90.1 83.6 87.6 88.2 73.9 66.0 68.6 79.8 0.4 0.1 0.3 16.2
Ollm-wiki-7B 74.3 84.5 67.2 81.4 77.0 65.2 83.3 53.4 57.3 59.3 79.0 0.1 0.0 0.2 0.1
Ollm-arxiv-7B 74.1 84.4 67.5 81.5 77.0 64.2 82.8 53.1 56.6 58.4 79.1 0.1 0.0 0.0 8.3
LLaMA3.1-8B 79.8 87.4 74.9 88.4 81.1 79.8 84.2 72.3 62.2 68.9 79.4 0.1 0.0 0.1 15.3
Ministral-8B 78.9 88.8 62.4 83.9 79.5 81.0 88.4 60.1 62.4 52.7 82.6 0.1 0.0 0.1 16.4
Internlm3-8B 83.1 90.9 72.0 88.9 82.4 88.5 90.5 73.8 67.2 62.9 79.7 0.2 0.0 0.4 12.0
Aya-8B 77.1 85.8 62.4 83.8 77.9 73.0 78.0 62.6 57.4 63.6 80.5 0.1 0.0 0.0 6.3
@@
Qwen2.5-14B 86.6 92.0 75.5 91.4 85.8 89.6 94.0 76.4 71.2 63.6 79.9 0.1 0.1 1.0 19.5
Mistral-22B 83.9 90.4 69.6 88.6 84.4 86.3 86.9 69.3 64.0 54.3 80.1 0.1 0.0 0.8 15.8
Qwen2.5-32B ❖ 88.0 90.6 81.9 91.2 87.2 89.7 95.5 76.8 72.4 68.4 80.0 1.6 0.1 1.5 20.3
QwQ-32B 82.2 89.6 77.1 88.9 81.5 84.0 92.5 70.8 60.6 63.4 79.4 1.1 0.2 1.0 18.0
Aya-32B 81.2 90.5 61.6 89.7 82.3 85.5 83.1 70.3 66.0 68.8 79.3 0.1 0.1 0.5 19.3
@@
LLaMA3.3-70B 88.0 94.1 76.8 91.8 90.0 91.9 92.9 76.8 70.9 64.2 80.0 0.1 0.0 0.7 20.2
Qwen2.5-72B ★ 89.1 92.6 84.3 92.6 89.4 92.1 93.4 77.5 75.6 68.4 79.9 0.1 0.0 1.0 21.6
Table 2: Main results (%) of 16 LLMs (grouped by size) under the zero-shot setting. ★ indicates the best-performing model overall, while ❖ denotes the best-performing model within its size category.
4.2.1 Performance of LLMs
@@
The two largest models, LLaMA3.3-70B and Qwen2.5-72B, consistently achieve the best performance. Notably, the Qwen architecture shows robust results at all sizes, outperforming other architectures of comparable scale. In contrast, Ollm, which is specifically trained for ontology construction, performs relatively poorly, likely due to its specialization in hierarchical generation rather than general understanding or reasoning.
@@
Model scale correlates strongly with performance, especially in understanding and reasoning. For instance, Qwen2.5-72B achieves top scores of 92.6 on U4, 93.4 on R2, and 21.6 F1-score on L5. Similarly, LLaMA3.3-70B scores 91.8 on U4, 92.9 on R2, and 20.2 on L5. This pattern is even more pronounced within the same architecture: across Qwen, Mistral, Aya, and LLaMA, larger models consistently perform better.
Ontological Understanding of LLMs
@@
Tasks U1 to U5 demonstrate that LLMs generally perform well on ontology understanding, particularly in recognizing hierarchical structures. This is reflected in the high accuracy on U2 (class relations) and U4 (instance classification), with scores ranging from 80% to 94%. In contrast, performance is less consistent on definition and property tasks. U1 (class definition), U3 (property domain), and U5 (instance definition) show notable gaps in certain models, for instance, Aya-8B scores only 77.1%, 62.4%, and 77.9% on these tasks.
Ontological Reasoning of LLMs
@@
Reasoning tasks present greater challenges for LLMs than understanding tasks. Task R1 (inferred relation reasoning) is a difficult variant of task U2 (class relation understanding), requiring reasoning to identify class relationships not explicitly defined. As expected, models generally perform 3-4 percentage points worse on R1 than on U2, with the most dramatic decrease observed in Ollm-arxiv-7B (from 84.4% to 64.2%). Similarly, R3 (instance class reasoning) functions as the reasoning-intensive counterpart to U4. Performance on R3 (60-70%) demonstrates a substantial decline compared to U4 (80-90%). These results indicate that when reasoning across multiple relationships is required, performance deteriorates significantly.
@@
Tasks involving logical expressions, such as R4 (SWRL-based reasoning) and R5 (description logic reasoning), are more difficult. Compared to natural language-based reasoning (R1, R2 and R3), model performance drops more dramatically when logical operators are involved, with scores ranging between 60% and 75%. Even the best-performing model, Qwen2.5-72B, achieves only 75.6% and 68.8% on these two tasks, respectively. This highlights a significant limitation of current LLMs: difficulty in executing precise symbolic reasoning over formally structured ontologies.
Ontological Learning of LLMs
@@
Although direct comparisons is limited due to different evaluation metrics, generation tasks are shown to be more challenging, comparing with understanding and reasoning tasks. In L1 (class definition generation), models struggle to generate definitions, as evidenced by low BERTScore (typically below 10). This poor performance likely stems from the challenges of domain-specific definition generation, which requires not only describing the target class but also distinguishing it from adjacent concepts (e.g. its superclasses). Unlike human domain experts who possess comprehensive ontological perspectives, LLMs struggle to make such fine-grained semantic distinctions within specialized domains.
@@
Similar limitations appear in structural construction tasks. In L2 (class hierarchy construction), L3 (property relation construction), and L4 (constraint construction), models frequently fail to produce syntactically valid triples in zero-shot settings. Performance improves modestly in two- and four-shot settings, but output quality remains low, often featuring ill-defined or hallucinated relations. For L5 (ontology alignment), models achieve slightly higher F1 scores in the range of 10-20%. But for practical ontology applications, the performance of the evaluated LLMs on all learning tasks remains substantially poor for reliable deployment.
4.2.2 Analysis
@@
Figure 5: The performance of Qwen2.5-7B, 72B (green and blue) and four domain-specific LLMs (yellow) across four domains.
In addition to the task‐based evaluation, Figure 5 compares the performance of two open-domain LLMs, and four domain‐specific LLMs (Finance-Chat-7B, OpenBioLLM-8B, BioMistral-7B, and SaulLM-7B). For the Sciences domain, we retained only the Biology‐related tasks. To simplify the computation of the mean scores for the learning task, we omitted the scores of L1.
As expected, the trends observed earlier hold as well: larger models consistently outperform smaller ones, and models struggle more on reasoning and learning than on understanding tasks. Comparing across domains, performance in Legal and the Sciences (Biology) lags behind the other two domains, with the gap most pronounced in Biology, which is likely a reflection of the greater complexity and specialized knowledge required.
@@
Does Concept Depth matter?
Concepts occupy different positions within the ontology and vary systematically in difficulty. As shown in Figure 6, model accuracy exhibits a consistent U-shaped pattern across all five models: performance is high at shallow depths (1–3), drops sharply at intermediate depths (4–8), and recovers beyond depth 10. This trend persists despite reduced sample sizes at greater depths, indicating that the recovery is not an artifact of data volume. We hypothesize that intermediate-depth concepts present a fundamental challenge: they are too specific to benefit from broad generalization yet too abstract to be directly memorized from pretraining data. The recovery at deeper levels suggests that highly specific, leaf-node concepts may contain sufficient distinctive features for easier classification. These results reveal a structural limitation in how LLMs process concepts at different levels of abstraction.
Figure 6: performance of five models with concept depth (0–20). The results aggregate tasks U1–U5 and R1–R3. Depths >> 20 are omitted due to scarce samples.
Does Few-shot Prompting help?
We evaluated 0-shot, 2-shot, and 4-shot prompting across five models. Few-shot examples yield modest improvements on Understanding and Reasoning tasks: Qwen-72B improves from 89.6% (0-shot) to 91.3% (4-shot). The most substantial gains occur on Learning tasks: Qwen-72B increases from 7.7% to 21.0%, and Qwen-32B from 5.7% to 19.1%. These improvements scale with model size, suggesting that few-shot effectiveness depends on both task complexity and model capacity. Notably, LLaMA-8B exhibits a slight performance decline on Reasoning with 4-shot, which we attribute to context length limitations in smaller models when processing longer prompts.
Figure 7: Performance of five models with 0, 2, and 4-shot prompting across three levels. Scores are aggregated within each task level. Full results are shown in Appendix E.
System Understanding Reasoning Learning
U1 U2 U3 U4 U5 R1 R2 R3 R4 R5 L1 L2 L3 L4 L5
Human1 55.0 60.0 65.0 65.0 65.0 75.0 70.0 75.0 65.0 70.0 3.0 20.1 51.0 49.0 51.7
Human2 65.0 60.0 60.0 75.0 75.0 80.0 70.0 65.0 60.0 60.0 2.0 30.1 48.3 61.0 53.0
Human3 60.0 60.0 62.5 70.0 70.0 77.5 70.0 70.0 62.5 65.0 2.5 25.2 49.5 55.0 52.5
Qwen-7B 83.1 90.6 77.6 90.1 83.6 87.6 88.2 73.9 66.0 68.6 5.6 0.4 0.1 0.3 16.2
Qwen-32B 88.0 90.6 81.9 91.2 87.2 89.7 95.5 76.8 72.4 68.4 5.7 1.6 0.1 1.5 20.3
Qwen-72B 89.1 92.6 84.3 92.6 89.4 92.1 93.4 77.5 75.6 68.4 6.0 0.1 0.0 1.0 21.6
Table 3: Comparison of humans and LLMs. Human results are computed over 30 randomly selected questions per task, without the help of any tools; LLM results are reported on the full dataset.
Does Chain of Thought (CoT) help?
We assess the impact of Chain-of-Thought prompting on performance. As shown in Figure 8, CoT prompting produces mixed results. On Understanding and Reasoning tasks, we observe modest to significant performance declines: Qwen2.5-7B drops from 83.1% to 80.9% on Understanding, while Qwen-72B shows larger decreases of 14.5pp and 15.6pp on Understanding and Reasoning, respectively. Conversely, Learning tasks demonstrate substantial improvements: Qwen-7B increases from 5.5% to 12.4%, Qwen-32B from 5.7% to 20.2%, and Qwen-72B from 6.0% to 18.5%.
These contrasting effects suggest that CoT’s impact is task-dependent. We propose two explanations: (1) Understanding and Reasoning tasks in OntoURL require ontology-specific inference patterns that differ fundamentally from the mathematical and commonsense reasoning prevalent in pretraining corpora, limiting CoT transferability; (2) Learning tasks benefit from CoT’s structured generation process, which helps models organize knowledge during content creation.
Figure 8: Performance impact of chain-of-thought prompting across five models, compared to standard prompting. Red indicates performance decreases, and blue indicates improvements. Full results are shown in Appendix E.
LLMs vs. Humans
To contextualize LLM capabilities, we conducted a human evaluation with three participants holding PhD-level expertise. As shown in Table 3, LLMs substantially outperform humans on Understanding tasks (average gap: +22.9pp), where formal ontological concepts in specialized domains (e.g., medical, science) pose significant challenges. The LLM advantage narrows on Reasoning tasks (average gap: +10.6pp), suggesting these tasks rely more on general logical reasoning—a skill less dependent on wide knowledge. Notably, humans achieve superior performance on Learning tasks (average gap: -31.6pp), though both groups struggle with definition generation (Task L1: humans 2.5%, LLMs 5.8%). Overall, the results suggest that LLMs excel at leveraging broad ontological knowledge, but face challenges in tasks requiring creative synthesis and generation.